# Multimodal Evaluation of Russian-language Architectures

### v2

@@ 3.4.1 Data Watermarking
  • Audio: We use AudioSeal roman2024proactivedetectionvoicecloning_watermarking for localized detection, which employs a neural, inaudible watermark.
@@ 3.4.2 Data Leakage Detection
We detect training-set data leakage using membership inference attacks shokri2017membership. Our approach extends the Semantic Membership Inference Attack (SMIA) mozaffari2024semanticmembershipinferenceattack_smia to MLLMs, calculating the loss for a data point by considering its text in conjunction with its paired image, video, or audio data. The Multimodal SMIA (MSMIA) method identifies leakage by comparing a model’s behavior on original examples versus their semantically perturbed "neighbors" (masked, removed, doubled, switched text tokens); models that have been trained on the data (models with data leak) are to exhibit systematically different loss patterns.
@@
Concretely, we train the MSMIA detector by comparing two versions of a model: the original and a version we fine-tune on candidate data (simulating a leak). The detector learns to distinguish between them by analyzing the differences in loss and text embeddings when processing original data points versus their perturbed neighbors. Once trained, this detector can be applied to any other model to output a probability that a specific data sample was part of that model’s training set. Overall results of the MSMIA detection capabilities are presented in Table 4. Following the original methodology, we evaluate detection performance using AUC-ROC. The table demonstrates relatively high scores, which means that the MSMIA method tends to be capable of detecting whether a model has been trained on some multimodal data samples or not, with high probability and a rather low false-positive rate. The details on the MSMIA training and the metrics are provided in Appendix D.
3.4.3 Legal Protection via Licensing
We introduce a bespoke legal framework, the MERA Multi License [^7] https://github.com/MERA-Evaluation/MERA_MULTIMODAL/blob/main/LICENSE, which explicitly permits the use of the benchmark data for research and non-commercial testing purposes, but strictly prohibits its incorporation into any model training process.
@@ 4.1 Model Baselines
See Appendix G.1 for the baseline details. We evaluate models in the same environments and scenarios by the procedure described in Section 3.3.4 and the submission procedure described in Section 3.3.5. We also provide examples of particular submissions (the model evaluated on part of the tasks of a modality).
@@ 4.2 Human Baselines
To estimate human-level performance across MERA Multi tasks, we compute human baseline (HB) values based on the aggregated results of crowd annotators. For datasets requiring additional domain expertise, we also establish expert HB obtained from qualified expert annotators. Annotation quality is ensured through honeypot tasks with automated correctness verification and post-hoc filtering of low-performing annotators. All crowd annotations are collected via the ABC Elementary platform [^8] https://app.elementary.center, which guarantees data anonymity, fair compensation, and ethical compliance. See Appendix G.2 for detailed methodology and cost breakdown .
@@ 5 Results
The leaderboard is designed in such a way that the more modalities the model covers, the higher the Total Score could be. The top performer, Qwen3-Omni-30B-A3B-Instruct, leads with a Total Score of 0.434, driven by its high Attempted Score (0.523) and rather high Coverage (0.828), showing strong image, audio, and video capabilities. Notably, the models from Qwen families obtain larger scores for image and video modalities (first 5 places of the overall leaderboard are taken by those models). GPT 4.1 still leads in image modality while having low Coverage (0.333), which leads to a lower Total Score (0.143 compared to 0.434 of the top performer). The main trend is defined by the metrics used: broader coverage leads to higher Total Score. Thus omni-models occupy the first places. But strong uni- or bimodal capabilities may gain advantage over middle-performing models with high Coverage (e.g. Qwen2.5-VL-72B-Instruct and Qwen2-VL-72B-Instruct with 0.257 and 0.251 Total Scores respectively).
This pattern is consistent across all modalities. In audio, the specialists from ultravox family tend to display poorer performance compared to omni-models like Qwen2.5-Omni-7B (0.268 vs 0.464 for Audio Total Score) even though ultravox models use other LLM’s from Mistral, Llama, Qwen families as backbones. Considering the video modality, vision models are usually trained with video-inputs or may slice the video into frames and use “regular” vision encoders for them, which explains why Qwen2-VL-72B-Instruct shows the best Video Total Score (0.54) while the models that specialize specifically on video modality like those from LLaVA-NeXT-Video family show poorer metrics.
@@
Tables with separate datasets metrics and analysis may be found in Appendix G.1.
Model
Total
Score
Attempted
Score
Coverage
Image
Total
Audio
Total
Video
Total
Qwen3-Omni-30B-A3B-Inst 0.434 0.523 0.828 0.423 0.531 0.357
Qwen2.5-Omni-7B 0.302 0.302 1.0 0.211 0.464 0.42
Qwen2.5-VL-72B-Inst 0.257 0.386 0.667 0.353 0.0 0.506
Qwen2-VL-72B-Inst 0.251 0.376 0.667 0.332 0.0 0.54
Qwen2.5-Omni-3B 0.244 0.244 1.0 0.172 0.377 0.332
MiniCPM-o-2_6 0.211 0.23 0.917 0.177 0.234 0.341
Qwen2.5-VL-7B-Inst 0.192 0.288 0.667 0.238 0.0 0.474
Qwen2-VL-7B-Inst 0.145 0.217 0.667 0.195 0.0 0.298
GPT 4.1 0.143 0.43 0.333 0.43 0.0 0.0
Qwen2.5-VL-3B-Inst 0.132 0.198 0.667 0.138 0.0 0.418
InternVL3-9B 0.128 0.192 0.667 0.169 0.0 0.276
Qwen3-VL-2B-Inst 0.121 0.182 0.667 0.124 0.0 0.396
Phi-4-multimodal-inst 0.097 0.145 0.667 0.183 0.042 0.0
ultravox-v0_4 0.089 0.268 0.333 0.0 0.268 0.0
ultravox-v0_5-llama-3_1-8b 0.089 0.266 0.333 0.0 0.266 0.0
ultravox-v0_4_1-llama-3_1-8b 0.088 0.264 0.333 0.0 0.264 0.0
audio-flamingo-3-hf 0.086 0.259 0.333 0.0 0.259 0.0
ultravox-v0_4_1-mistral-nemo 0.086 0.257 0.333 0.0 0.257 0.0
ultravox-v0_6-llama-3_1-8b 0.082 0.247 0.333 0.0 0.247 0.0
llava-next-110b-hf 0.078 0.235 0.333 0.235 0.0 0.0
Phi-3.5-vision-inst 0.075 0.226 0.333 0.226 0.0 0.0
Qwen2-Audio-7B-Inst 0.074 0.223 0.333 0.0 0.223 0.0
SmolVLM-Inst 0.063 0.189 0.333 0.189 0.0 0.0
gemma-3-27b-it 0.049 0.146 0.333 0.146 0.0 0.0
granite-vision-3.3-2b 0.047 0.142 0.333 0.142 0.0 0.0
deepseek-vl2-small 0.044 0.162 0.273 0.132 0.0 0.0
@@ Conclusion
The rapid progress of generative AI has introduced new challenges for evaluating models in multimodal contexts. We present MERA Multi, the first comprehensive framework for transparent and culturally grounded multimodal evaluation in Russian. The benchmark encompasses 18 tasks across four modalities (default text, image, audio, and video), covering diverse domains and scenarios. It systematizes diverse multimodal abilities via a proposed taxonomy and evaluates them through methodologically verified prompts and metrics. We also provide a standardized code base [^9] https://github.com/MERA-Evaluation/MERA_MULTIMODAL that guarantees reproducibility and a submission platform [^10] https://mera.a-ai.ru/ru/multi offering automated evaluation, scoring, and open leaderboards.
@@
For the creation of novel datasets, the work of human annotators is used. We state that their work was adequately paid or compensated (see Section G.2 for the details).
@@ Acknowledgments
We are profoundly grateful to the following individuals for their dedication and hard work: Yaroslav Grebnyak, Anna Kostikova, Aleksandr Sautin, Artem Goryainov, Aleksandra Rak, Artem Goryainov, Albina Akhmetgareeva, Igor Churin, Leonid Sinev, Yulia Lazareva, Ksenia Biryukova, Jamilya Erkenova, Valentina Khlebutina, Mariia Slabunova, Sergei Markov and to many others whom we may have inadvertently missed, but who supported us with their ideas, collaborated on dataset creation, validated results, and helped organize the workflow — your contributions are sincerely appreciated.
— The MERA Team
@@
Appendix A Multimodal Benchmarks Comparison
Benchmark Modalities # Tasks Primary Focus Language Cultural Focus
MultiBench liang2021multibench 10 (Text, Image, Audio, Video, …) 20 General-purpose representation EN General
MMBench liu2024mmbench Image, Text 1 Fine-grained visual reasoning EN, ZH General
SUPERB yang2021superb Speech, Audio 10 Speech processing EN General
STAR wu2024star Video, Text 4 VideoQA, temporal reasoning EN General
InfiniBench ataallah2024infinibench Video, Text 8 Long-context video understanding EN General
Video-MME fu2025video Video, Text 1 Fine-grained video analysis EN General
General-Bench fei2025path Text, Image, Video, Audio, Tabular 700+ Large-scale ability coverage EN General
OmniDialog razzhigaev2024omnidialog Text, Image, Audio 8 Multimodal dialogue EN General
MMMU yue2023mmmu Image, Text 30 Expert-level reasoning EN General
SEED-Bench li2023seed Image, Video, Text 12 Generative comprehension EN General
MERA Multi Text, Image, Audio, Video 18 Comprehensive understanding RU General + Russian
Table 6 presents the comparison of major multimodal evaluation benchmarks with MERA Multi.
Appendix B Dataset Description
B.1 AQUARIA
@@
The methodology for evaluating large audio-language models (LALMs), as well as the models themselves, is a fairly recent area of research. Compared to benchmarks in the vision-language domain, there are significantly fewer comprehensive benchmarks available for evaluating audio-language models. Examples of such benchmarks include AIR-Bench yang2024airbenchbenchmarkinglargeaudiolanguage, AudioBench wang2025audiobenchuniversalbenchmarkaudio, and MMAU sakshi2024mmaumassivemultitaskaudio. Audio understanding tasks are generally classified into three categories: speech analysis, non-verbal signal analysis, and music analysis.
@@ B.5 RealVQA
B.6 ruCLEVR
RuCLEVR is a Visual Question Answering (VQA) dataset inspired by the CLEVR johnson2016clevrdiagnosticdatasetcompositional methodology and adapted for the Russian language.
@@
The dataset addresses the classic foundational Visual Question Answering (VQA) task, similar to English datasets such as VQA balanced_vqa_v2. Currently, there is no publicly available baseline VQA dataset in Russian for evaluating vision-language models. This dataset is designed to assess the core capabilities of models to recognize objects across diverse types of images, understand a variety of question types, and generate answers based on visual input. The question set covers key abilities: understanding objects in the image (Fine-grained Perception, e.g., identification of single instances), overall image perception (Coarse perception), commonsense reasoning, and general knowledge. Images are sourced from public datasets, including COCO lin2015microsoftcococommonobjects and English-language VQA v2161616https://hf.co/datasets/pingzhili/vqa_v2, which should be considered as limitation when interpreting evaluation results. There is a possibility of indirect data leakage through the images in model training data.
@@
To construct the dataset, images were sourced from the English-language VQA v2 dataset (which includes data from the COCO lin2015microsoftcococommonobjects dataset). Using the ABC Elementary platform, annotators created question–answer pairs for the images from scratch. Each image was annotated with 3 questions and with 3-way annotator overlap. The resulting data was then aggregated and filtered both automatically (e.g., removal of overly long answers, typos, formatting issues) and manually. The binary question data was class-balanced.
@@
B.8 ruEnvAQA
ruEnvAQA is a dataset of multiple-choice and binary-choice questions in Russian. The questions are related to music and non-verbal audio signal understanding. The dataset is based on questions from English-language datasets Clotho-AQA lipping2022clothoaqacrowdsourceddatasetaudio and MUSIC-AVQA li2022learninganswerquestionsdynamic. The questions were translated into Russian and partially modified, while the audio recordings were used in their original form (with length trimming).
@@
Compared to the vision-language domain, there are fewer large benchmarks that combine diverse tasks for the evaluation of LALM skills. Examples of such benchmarks include AIR-Bench yang2024airbenchbenchmarkinglargeaudiolanguage, AudioBench wang2025audiobenchuniversalbenchmarkaudio, and MMAU sakshi2024mmaumassivemultitaskaudio. Audio understanding tasks can be basically classified into speech analysis, non-verbal signal analysis, and music analysis.
@@
The Clotho-AQA lipping2022clothoaqacrowdsourceddatasetaudio dataset contains questions about audio with non-verbal signals and minor speech elements, with questions focusing only on non-verbal signals and occasionally on external characteristics of speech, such as volume or speaker gender.
@@
The MUSIC-AVQA li2022learninganswerquestionsdynamic dataset consists of video recordings of musical performances and three groups of questions:
@@
B.12 ruNaturalScienceVQA
NaturalScienceQA is a multimodal question-answering dataset on natural sciences with basic questions from the school curriculum, based on the English dataset ScienceQA lu2022learn. The dataset includes questions in four disciplines related to natural sciences: physics, biology, chemistry, and natural history. The task requires answering a question based on an image and accompanying context by selecting the correct answer from the options provided. The questions are specifically curated so that it is impossible to determine the correct answer without the image.
@@
NaturalScienceQA was created based on the English ScienceQA lu2022learn dataset, a question-answering dataset covering a wide range of scientific disciplines. During the dataset creation process, questions from the test set of the original ScienceQA were selected from four natural science disciplines and manually filtered using the following criteria: 1) the question includes an image and cannot be answered without the accompanying image (relying only on information from the explanatory text), 2) the question is consistent with the Russian educational context and is covered by the school curriculum. Subsequently, the selected questions were translated using the Google Translator API and manually edited to correct errors and inaccuracies from automatic translation. Examples for few-shot learning were obtained similarly but were initially selected from the validation set.
B.13 ruSLUn
RuSLUn (Russian Spoken Language UNderstanding dataset) is a Russian-language dataset for spoken language understanding, designed following the principles of the English SLURP bastianelli2020slurpspokenlanguageunderstanding and the multilingual xSID vandergoot2021maskedlanguagemodelingtranslation, but with consideration for the cultural and linguistic specifics of Russia. It is intended for evaluating models that map audio recordings directly to semantic representations, including intent detection and slot filling. RuSLUn contains a variety of spoken commands and queries that are typical for Russian users and contexts. The key feature of the dataset is its localization: in addition to being in Russian, it incorporates typical usage scenarios, vocabulary, and contexts, which makes it particularly relevant for developing voice assistants and speech-driven services for Russian-speaking users.
@@
The dataset was created in two stages: first, text queries were generated and annotated with intents and slots, then, these queries were recorded as audio. The annotation scheme was based on the cross-lingual xSID vandergoot2021maskedlanguagemodelingtranslation dataset, which includes 16 intent types and 33 slot types. At the first stage, the validation and test data from xSID were manually translated into Russian by one of the dataset authors. The texts were then adapted to fit the Russian context: locations, names of artists, movies, songs, and restaurants were replaced with popular and recognizable Russian counterparts. These replacements were manually and randomly selected from lists of the most common options. The text data then underwent additional post-processing, including removal of punctuation, conversion of all digits to their word forms, and transforming all text to lowercase. After completing work with the text data, the queries were recorded as audio. Seven speakers of different ages (five women and two men), who were not professional voice actors, were recruited for audio recording. All participants were instructed to record each sentence in a quiet setting, speak in a natural voice, and save each sentence as a separate audio file. Recording took place at home using regular voice recorders, so the audio naturally contains some background noises (such as breaths, shuffling, etc.). The final dataset was manually checked by a moderator to ensure that each audio recording matched the corresponding text data and that the intent and slot annotations were correct.
B.14 ruTiE-Audio
@@
B.16 SchoolScienceVQA
SchoolScienceVQA is a Russian-language multimodal dataset inspired by ScienceQA lu2022learn. It evaluates the reasoning capabilities of AI models in a multimodal setting using multiple-choice questions across scientific subjects such as physics, biology, chemistry, economics, history, and earth science. Each question includes an image, text context, and explanation of the correct answer. These components provide a basis for assessing reasoning chains.

### v3

@@ 3.4.1 Data Watermarking
  • Audio: We use AudioSeal Roman et al. (2024) for localized detection, which employs a neural, inaudible watermark.
@@ 3.4.2 Data Leakage Detection
We detect training-set data leakage using membership inference attacks Shokri et al. (2017). Our approach Emelyanov et al. (2025) extends the Semantic Membership Inference Attack (SMIA) Mozaffari and Marathe (2024) to MLLMs, calculating the loss for a data point by considering its text in conjunction with its paired image, video, or audio data. The Multimodal SMIA (MSMIA) method identifies leakage by comparing a model’s behavior on original examples versus their semantically perturbed "neighbors" (masked, removed, doubled, switched text tokens); models that have been trained on the data (models with data leak) are to exhibit systematically different loss patterns.
@@
Concretely, we train the MSMIA detector by comparing two versions of a model: the original and a version we fine-tune on candidate data (simulating a leak). The detector learns to distinguish between them by analyzing the differences in loss and text embeddings when processing original data points versus their perturbed neighbors. Once trained, this detector can be applied to any other model to output a probability that a specific data sample was part of that model’s training set. Overall results of the MSMIA detection capabilities are presented in Table 5. Following the original methodology, we evaluate detection performance using AUC-ROC. The table demonstrates relatively high scores, which means that the MSMIA method tends to be capable of detecting whether a model has been trained on some multimodal data samples or not, with high probability and a rather low false-positive rate. The details on the MSMIA training and the metrics are provided in Appendix C.
@@ 4.1 Model Baselines
See Appendix F.1 for the baseline details. We evaluate models in the same environments and scenarios by the procedure described in Section 3.3.4 and the submission procedure described in Section 3.3.5. We also provide examples of particular submissions (the model evaluated on part of the tasks of a modality).
@@ 4.2 Human Baselines
To estimate human-level performance across MERA Multi tasks, we compute human baseline (HB) values based on the aggregated results of crowd annotators. For datasets requiring additional domain expertise, we also establish expert HB obtained from qualified expert annotators. Annotation quality is ensured through honeypot tasks with automated correctness verification and post-hoc filtering of low-performing annotators. All crowd annotations are collected via the ABC Elementary platform [^9] https://app.elementary.center, which guarantees data anonymity, fair compensation, and ethical compliance. See Appendix F.2 for detailed methodology and cost breakdown .
Model Total Score Attempted Score Coverage Image Total Audio Total Video Total
Qwen3-Omni-30B-A3B-Inst 0.500 0.563 0.889 0.554 0.561 0.410
Qwen2.5-Omni-7B 0.317 0.317 1.000 0.226 0.474 0.442
Qwen2.5-VL-72B-Inst 0.302 0.453 0.667 0.406 0.000 0.625
MiniCPM-o-2_6 0.255 0.255 1.000 0.182 0.369 0.373
Qwen2-VL-72B-Inst 0.254 0.381 0.667 0.333 0.000 0.557
Qwen2.5-Omni-3B 0.251 0.251 1.000 0.180 0.380 0.337
Qwen2.5-VL-7B-Inst 0.209 0.313 0.667 0.256 0.000 0.523
GPT 4.1 0.159 0.478 0.333 0.478 0.000 0.000
Qwen2-VL-7B-Inst 0.145 0.218 0.667 0.195 0.000 0.301
Qwen2.5-VL-3B-Inst 0.136 0.203 0.667 0.142 0.000 0.427
InternVL3-9B 0.135 0.203 0.667 0.172 0.000 0.316
Qwen3-VL-2B-Inst 0.125 0.187 0.667 0.125 0.000 0.416
ultravox-v0_5-llama-3_1-8b 0.104 0.311 0.333 0.000 0.311 0.000
ultravox-v0_4_1-llama-3_1-8b 0.102 0.307 0.333 0.000 0.307 0.000
ultravox-v0_4 0.101 0.304 0.333 0.000 0.304 0.000
ultravox-v0_6-llama-3_1-8b 0.100 0.300 0.333 0.000 0.300 0.000
Phi-4-multimodal-instruct 0.098 0.147 0.667 0.185 0.042 0.000
ultravox-v0_4_1-mistral-nemo 0.088 0.265 0.333 0.000 0.265 0.000
audio-flamingo-3-hf 0.086 0.259 0.333 0.000 0.259 0.000
llava-next-110b-hf 0.079 0.236 0.333 0.236 0.000 0.000
Phi-3.5-vision-inst 0.076 0.228 0.333 0.228 0.000 0.000
Qwen2-Audio-7B-Inst 0.074 0.223 0.333 0.000 0.223 0.000
SmolVLM-Inst 0.064 0.192 0.333 0.192 0.000 0.000
gemma-3-27b-it 0.050 0.151 0.333 0.151 0.000 0.000
granite-vision-3.3-2b 0.048 0.143 0.333 0.143 0.000 0.000
deepseek-vl2-small 0.044 0.163 0.273 0.133 0.000 0.000
@@ 5 Results
The leaderboard is designed in such a way that the more modalities the model covers, the higher the Total Score could be. The top performer, Qwen3-Omni-30B-A3B-Instruct, leads with a Total Score of 0.5, driven by its high Attempted Score (0.563) and rather high Coverage (0.889), showing strong image, audio, and video capabilities. Notably, the models from Qwen families obtain larger scores for image and video modalities (first 3 places of the overall leaderboard are taken by those models). GPT 4.1 still leads in image modality while having low Coverage (0.333), which leads to a lower Total Score (0.159 compared to 0.5 of the top performer). The main trend is defined by the metrics used: broader coverage leads to higher Total Score. Thus omni-models occupy the first places. But strong uni- or bimodal capabilities may gain advantage over middle-performing models with high Coverage (e.g. Qwen2.5-VL-72B-Instruct and Qwen2-VL-72B-Instruct with 0.302 and 0.254 Total Scores respectively).
This pattern is consistent across all modalities. In audio, the specialists from ultravox family tend to display poorer performance compared to omni-models like Qwen2.5-Omni-7B (0.311 vs 0.474 for Audio Total Score) even though ultravox models use other LLM’s from Mistral, Llama, Qwen families as backbones. Considering the video modality, vision models are usually trained with video-inputs or may slice the video into frames and use “regular” vision encoders for them, which explains why Qwen2-VL-72B-Instruct shows the best Video Total Score (0.625) while the models that specialize specifically on video modality like those from LLaVA-NeXT-Video family show poorer metrics.
@@
Tables with separate datasets metrics and analysis may be found in Appendix F.1.
@@ Conclusion
The rapid progress of generative AI has introduced new challenges for evaluating models in multimodal contexts. We present MERA Multi, the first comprehensive framework for transparent and culturally grounded multimodal evaluation in Russian. The benchmark encompasses 18 tasks across four modalities (default text, image, audio, and video), covering diverse domains and scenarios. It systematizes diverse multimodal abilities via a proposed taxonomy and evaluates them through methodologically verified prompts and metrics. We also provide a standardized code base [^10] https://github.com/MERA-Evaluation/MERA_MULTIMODAL that guarantees reproducibility and a submission platform offering automated evaluation, scoring, and open leaderboards.
@@
For the creation of novel datasets, the work of human annotators is used. We state that their work was adequately paid or compensated (see Section F.2 for the details).
@@ Acknowledgments
We are profoundly grateful to the following individuals for their dedication and hard work: Yaroslav Grebnyak, Anna Kostikova, Aleksandr Sautin, Artem Goryainov, Aleksandra Rak, Artem Goryainov, Albina Akhmetgareeva, Igor Churin, Leonid Sinev, Yulia Lazareva, Ksenia Biryukova, Jamilya Erkenova, Valentina Khlebutina, Maria Slabunova, Sergei Markov and to many others whom we may have inadvertently missed, but who supported us with their ideas, collaborated on dataset creation, validated results, and helped organize the workflow — your contributions are sincerely appreciated.
We also thank Zaryana Damashova and Ekaterina Artemova for their contributions to the creation of the RuSLUn dataset.
The work of Alexander Panchenko was supported by the RSF project 25-71-30008 “Laboratory for reliable, adaptive, and trustworthy Artificial Intelligence”.
@@
Appendix A Dataset Description
A.1 AQUARIA
@@
The methodology for evaluating large audio-language models (LALMs), as well as the models themselves, is a fairly recent area of research. Compared to benchmarks in the vision-language domain, there are significantly fewer comprehensive benchmarks available for evaluating audio-language models. Examples of such benchmarks include AIR-Bench Yang et al. (2024), AudioBench Wang et al. (2025a), and MMAU Sakshi et al. (2025). Audio understanding tasks are generally classified into three categories: speech analysis, non-verbal signal analysis, and music analysis.
@@ A.5 RealVQA
A.6 ruCLEVR
RuCLEVR is a Visual Question Answering (VQA) dataset inspired by the CLEVR Johnson et al. (2017) methodology and adapted for the Russian language.
@@
The dataset addresses the classic foundational Visual Question Answering (VQA) task, similar to English datasets such as VQA Goyal et al. (2017). Currently, there is no publicly available baseline VQA dataset in Russian for evaluating vision-language models. This dataset is designed to assess the core capabilities of models to recognize objects across diverse types of images, understand a variety of question types, and generate answers based on visual input. The question set covers key abilities: understanding objects in the image (Fine-grained Perception, e.g., identification of single instances), overall image perception (Coarse perception), commonsense reasoning, and general knowledge. Images are sourced from public datasets, including COCO Lin et al. (2014) and English-language VQA v2161616https://hf.co/datasets/pingzhili/vqa_v2, which should be considered as limitation when interpreting evaluation results. There is a possibility of indirect data leakage through the images in model training data.
@@
To construct the dataset, images were sourced from the English-language VQA v2 dataset (which includes data from the COCO Lin et al. (2014) dataset). Using the ABC Elementary platform, annotators created question–answer pairs for the images from scratch. Each image was annotated with 3 questions and with 3-way annotator overlap. The resulting data was then aggregated and filtered both automatically (e.g., removal of overly long answers, typos, formatting issues) and manually. The binary question data was class-balanced.
@@
A.8 ruEnvAQA
ruEnvAQA is a dataset of multiple-choice and binary-choice questions in Russian. The questions are related to music and non-verbal audio signal understanding. The dataset is based on questions from English-language datasets Clotho-AQA Lipping et al. (2022) and MUSIC-AVQA Li et al. (2022). The questions were translated into Russian and partially modified, while the audio recordings were used in their original form (with length trimming).
@@
Compared to the vision-language domain, there are fewer large benchmarks that combine diverse tasks for the evaluation of LALM skills. Examples of such benchmarks include AIR-Bench Yang et al. (2024), AudioBench Wang et al. (2025a), and MMAU Sakshi et al. (2025). Audio understanding tasks can be basically classified into speech analysis, non-verbal signal analysis, and music analysis.
@@
The Clotho-AQA Lipping et al. (2022) dataset contains questions about audio with non-verbal signals and minor speech elements, with questions focusing only on non-verbal signals and occasionally on external characteristics of speech, such as volume or speaker gender.
@@
The MUSIC-AVQA Li et al. (2022) dataset consists of video recordings of musical performances and three groups of questions:
@@
A.12 ruNaturalScienceVQA
NaturalScienceQA is a multimodal question-answering dataset on natural sciences with basic questions from the school curriculum, based on the English dataset ScienceQA Lu et al. (2022). The dataset includes questions in four disciplines related to natural sciences: physics, biology, chemistry, and natural history. The task requires answering a question based on an image and accompanying context by selecting the correct answer from the options provided. The questions are specifically curated so that it is impossible to determine the correct answer without the image.
@@
NaturalScienceQA was created based on the English ScienceQA Lu et al. (2022) dataset, a question-answering dataset covering a wide range of scientific disciplines. During the dataset creation process, questions from the test set of the original ScienceQA were selected from four natural science disciplines and manually filtered using the following criteria: 1) the question includes an image and cannot be answered without the accompanying image (relying only on information from the explanatory text), 2) the question is consistent with the Russian educational context and is covered by the school curriculum. Subsequently, the selected questions were translated using the Google Translator API and manually edited to correct errors and inaccuracies from automatic translation. Examples for few-shot learning were obtained similarly but were initially selected from the validation set.
A.13 ruSLUn
RuSLUn (Russian Spoken Language UNderstanding dataset) is a Russian-language dataset for spoken language understanding, designed following the principles of the English SLURP Bastianelli et al. (2020) and the multilingual xSID van der Goot et al. (2021), but with consideration for the cultural and linguistic specifics of Russia. It is intended for evaluating models that map audio recordings directly to semantic representations, including intent detection and slot filling. RuSLUn contains a variety of spoken commands and queries that are typical for Russian users and contexts. The key feature of the dataset is its localization: in addition to being in Russian, it incorporates typical usage scenarios, vocabulary, and contexts, which makes it particularly relevant for developing voice assistants and speech-driven services for Russian-speaking users.
@@
The dataset was created in two stages: first, text queries were generated and annotated with intents and slots, then, these queries were recorded as audio. The annotation scheme was based on the cross-lingual xSID van der Goot et al. (2021) dataset, which includes 16 intent types and 33 slot types. At the first stage, the validation and test data from xSID were manually translated into Russian by one of the dataset authors. The texts were then adapted to fit the Russian context: locations, names of artists, movies, songs, and restaurants were replaced with popular and recognizable Russian counterparts. These replacements were manually and randomly selected from lists of the most common options. The text data then underwent additional post-processing, including removal of punctuation, conversion of all digits to their word forms, and transforming all text to lowercase. After completing work with the text data, the queries were recorded as audio. Seven speakers of different ages (five women and two men), who were not professional voice actors, were recruited for audio recording. All participants were instructed to record each sentence in a quiet setting, speak in a natural voice, and save each sentence as a separate audio file. Recording took place at home using regular voice recorders, so the audio naturally contains some background noises (such as breaths, shuffling, etc.). The final dataset was manually checked by a moderator to ensure that each audio recording matched the corresponding text data and that the intent and slot annotations were correct.
A.14 ruTiE-Audio
@@
A.16 SchoolScienceVQA
SchoolScienceVQA is a Russian-language multimodal dataset inspired by ScienceQA Lu et al. (2022). It evaluates the reasoning capabilities of AI models in a multimodal setting using multiple-choice questions across scientific subjects such as physics, biology, chemistry, economics, history, and earth science. Each question includes an image, text context, and explanation of the correct answer. These components provide a basis for assessing reasoning chains.